OpenStudy (anonymous):

Hey I really, really need help with this problem: Travis is riding the Ferris wheel at the amusement park. His height can be modeled by the equation H(t) = 22 cos (pi/13) t + 28, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Travis before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Travis begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Travis' height above the ground reaches a minimum?

OpenStudy (anonymous):

@ganeshie8 @kc_kennylau @ranga @shamil98 @AriPotta @ganeshie8 @TuringTest @thomaster

OpenStudy (anonymous):

@Tazmann @ganeshie8 @TuringTest @thomaster

OpenStudy (anonymous):

Part 1, what is the time before the right begins? And what does the formula give as an answer if you fill in that number for t?

OpenStudy (anonymous):

Not sure, how do i find this?

OpenStudy (anonymous):

I assume t=0, since must formula start from t=0 (doesn't really make much sense to start from anywhere else).

OpenStudy (anonymous):

H(0) = 22 cos (pi/13) 28 ?

OpenStudy (anonymous):

H(0) = 22*cos(pi/13)*0 + 28

OpenStudy (anonymous):

okay so now what?

OpenStudy (anonymous):

By the, way, are you sure that the formula you gave at the first post isn't more like H(t) = 22 cos (pi/13t) + 28 or something? because otherwise you have a straight line, that doesn't make sense.

OpenStudy (anonymous):

\[\large 22 \cos (\frac{ \pi }{ 13 }t)+28\] by any chance?

OpenStudy (anonymous):

no thats the formula given. the first one

OpenStudy (anonymous):

That doesn't make sense, the t should be part of the cosine function or it would give a straight line, and then the questions don't really make sense.

OpenStudy (anonymous):

i"ll attach a pic of the equation

OpenStudy (anonymous):

there

OpenStudy (anonymous):

... alright then.

OpenStudy (anonymous):

In that case, the answer to part one is 28

OpenStudy (anonymous):

The answer to part 2 is forever, since according to that equation, the ferris wheel does not turn at all.

OpenStudy (anonymous):

And part 3 is not even possible to answer, since the top would be the edge of the universe and would even not be anywhere near earth.

OpenStudy (anonymous):

whats you reasoning for the first answer? how did you get 28?

OpenStudy (anonymous):

The 0 in the equation cancels out everything except the +28. Anything multiplied by 0 is zero.

OpenStudy (anonymous):

That is assuming the ride starts at t=0

OpenStudy (anonymous):

Okay thanks a lot. Can you help me with one more? i think its kinda easy but i"m not sure.

OpenStudy (anonymous):

I'l close this an open a new question

OpenStudy (anonymous):

I'm pretty sure the question was written incorrectly since it came from FLVS and they are known to have a lot of mistakes in their math curriculum...it's pretty pathetic